# What MLLMs Learn about When they Learn about Multimodal Reasoning: Perception, Reasoning, or their Integration?

### v2

@@ 3.1 Validating MathLens as a multimodal reasoning benchmark
To test whether MathLens captures patterns consistent with prior benchmarks, we compare it to MathVista (Lu et al. (2024)), MathVerse (Zhang et al. (2024)), and MathVision (Wang et al. (2024b)) by correlating model accuracies across eight models (Figure 5, left; full results in Appendix). MathLens shows strong Spearman’s $\rho$ correlation with MathVista ($\rho=0.83$) and MathVerse ($\rho=0.86$), confirming alignment with established benchmarks, while correlation with MathVision is weaker but still positive ($\rho=0.67$). Importantly, our goal is not a new downstream benchmark but a controlled, decomposition-focused resource; high correlations therefore underscore its consistency.
@@ 3.2 How much of the gain is explained by improvement in textual reasoning?
Figure 5: (left) Correlation of MathLens with popular benchmarks. MathLens shows high correlation with standard multimodal reasoning benchmarks. (right) Performance gains by input modality. Bars show percentage point shifts from finetuning for text versus diagram inputs. Visual gains exceed textual ones when models are primed with strong reasoning (textual SFT).
@@
Figure 5 (right) shows that textual SFT mainly improves textual reasoning, while multimodal RL applied afterward yields larger multimodal gains by shifting its effect to perception and integration. Direct multimodal RL without textual SFT gives only modest improvements in both modalities, and poor-quality multimodal SFT degrades multimodal performance more than textual reasoning, with RL only partially recovering the gap.
Finding 1: Multimodal RL impact varies with textual reasoning strength. With strong textual SFT it mainly boosts perception; without it, it modestly improves both modalities (fig. 5).
@@ 3.3 How much of the gain is explained by improvement in perception?
Figure 6: (left) Perception performance shifts from finetuning in percentage points. Except for multimodal SFT, all methods contribute to increased perception capacity. (right) Sample distribution over reflection count. Reflection is more frequent in false→true cases than in true→true cases, showing that perception gains from textual SFT arise partially from reflective reasoning.
Next, we evaluate how multimodal reasoning training affects low-level perception. For each problem $Q_{k}$, we use the perception probes $Q^{\mathrm{perc}}_{k}={q^{\mathrm{perc}}_{k,1},\ldots,q^{\mathrm{perc}}_{k,j}}$. Figure 6 (left) shows that all RL models contribute to better perception required for geometry problem solving. Hence, we conclude that perception is elicited even by training with correctness reward signals on the downstream problems.
@@
Finding 2: Textual SFT also improves perception by fostering reflective reasoning (fig. 6).
@@ 3.4 Error type analysis
MathLens’s annotation designs enable systemic categorization of model outputs into the following error categories: (1) Perception & Reasoning, failures on both perception probes and textual reasoning; (2) Perception, failure on perception probes but correct reasoning from text; (3) Reasoning, correct perception probes but failed text-based reasoning; (4) Integration, correct perception and text in isolation but failure on the combined multimodal task. Correct categories: (5) Trivial, solvable from text alone; and (6) Rest, all other correct cases.
Figure 1 (see full results in Figure 15 of Appendix) shows that RL primarily reduces perception and reasoning errors on the same problems, indicating correlated gains. Yet many of the reductions manifest as Integration errors, suggesting that finetuning often shifts them into coordination failures.
Finding 3: RL improves perception and reasoning in a correlated manner, and unveils integration as the remaining source of error once the other capacities improve (fig. 1).
@@ 3.5 Does multi-modal finetuning effect model’s robustness to visual inputs?
We extend the visual familiarity analysis from Section 3.1 by testing robustness under controlled diagram variations. Instead of pixel-level augmentations (e.g., blurring), we apply semantic modifications to the geometric specification, isolating structural variation without introducing low-level artifacts (see Appendix C).
@@
Figure 7 shows that while most training methods are stable under familiar diagrams, multimodal SFT suffers a sharp drop on modified diagrams. This indicates that similar downstream accuracy on public benchmarks can hide reliance on visual familiarity.
Figure 7: Finetuning shifts by diagram type. Bars show percentage point changes for existing versus out-of-distribution rendered diagrams.
@@
To quantify robustness, we use Consistency Rate (CR) (Zhao et al. (2024)), the expected agreement of predictions across perturbations of the same diagram:
$CR$ $=\mathbb{E}_{Q^{\prime}_{k}\sim Q^{\prime}}\Big[\;\mathbb{E}_{\begin{subarray}{c}\bar{C}_{k,i},\,\bar{C}_{k,j}\sim\bar{C}_{k}\\ i\neq j\end{subarray}}\mathbf{1}\!\left[M(Q^{\prime}_{k},\bar{C}_{k,i})=M(Q^{\prime}_{k},\bar{C}_{k,j})\right]\;\Big],$ (1)
@@
Figure 8 shows that accuracy on base diagrams is generally higher than on modified ones, indicating vulnerability to semantic perturbations. Consistency increases after multimodal RL, suggesting robustness to variation. By contrast, multimodal SFT lowers consistency.
Finding 4: Multimodal RL improves visual consistency under structural variations (fig. 8).
@@ 3.6 Which perception skills benefit from multimodal reasoning training?
We decompose perception probe performance (Section 3.3) by relation type (full details in Section E.4). This analysis reveals uneven gains, as shown in Figure 9. Relation types such as cocircular, parallel, and collinear improve consistently, as they rely on simple primitives. same_angle, val_angle, triangle, and quadrilateral improve only with textual SFT plus multimodal RL, since they require multi-constraint reasoning or symbol–geometry links. By contrast, same_length, val_length, and perpendicular remain difficult, as their cues are spatially offset or visually ambiguous.
Finding 5: Perception gains are uneven. Direct geometric cues improve reliably, while tasks relying on symbolic marks or distant annotations remain difficult (fig. 9).
@@ 4 Conclusion
We introduced MathLens, a controlled benchmark that disentangles perception, reasoning, integration, and robustness in multimodal reasoning. Our findings show that reinforcement learning primarily boosts perception, with stronger gains when supported by textual supervision, while textual SFT indirectly strengthens perception through reflective reasoning. Reasoning improves in tandem with perception under RL but does not exhibit distinct additional gains, leaving integration as the least improved capacity and the dominant failure mode once other skills advance. Robustness further diverges across strategies, as RL enhances consistency under diagram variation, whereas multimodal SFT reduces it through overfitting.
Looking ahead, our results motivate future architectures and training strategies that explicitly target integration, for example by introducing auxiliary pretext objectives for RL that enforce cross-modal grounding, or by structuring training data to better capture causal correspondences between perceptual details and reasoning trajectories. In parallel, scaling atomic perception probes into auxiliary supervision offers a promising direction for directly improving perceptual capacity.
5 Acknowledgements
We sincerely thank Vidhisha Balachandran, Shivam Garg, Jyoti Aneja, and Tyler LaBonte for their invaluable conversations and feedback throughout this project.
Ethics statement
MathLens consists solely of mathematical problems and does not involve human subjects or sensitive data. MathLens-General is constructed as an extension of existing benchmarks, and ethical considerations are therefore inherited from those sources. All human annotations were performed directly by the authors. As the benchmark is derived from publicly available resources, it does not raise additional privacy or copyright concerns beyond those already addressed in the original sources. As the content is limited to mathematical problems, concerns of bias, fairness, or harmful applications are not applicable.
Reproducibility statement
All experiments are conducted using existing models without additional training. The complete list of models is provided in Table 1, with hyperparameter configurations in Appendix D. Results are reported under deterministic decoding, aside from a small subset of models requiring random sampling to mitigate repetition. Minor nondeterminism from computation kernels, common across current LLM decoding environments (both local and API-based), is generally not treated as a controlled factor (He & Lab, 2025). As API services may evolve over time with undocumented changes, our main experiments focus on open-weight models, with API model results reported as auxiliary reference. The full dataset and model outputs will be released for reproducibility.
@@ Appendix A Related Work
Multimodal reasoning is a compositional process that requires the integration of perception and abstract reasoning (Li et al. (2025)). Inspired by textual reasoning benchmarks, recent multimodal reasoning datasets focus on verifiable domains such as mathematics (Lu et al. (2024); Wang et al. (2024b); Zhang et al. (2024)), scientific diagrams (Yue et al. (2024a); Hao et al. (2025)), and charts (Wang et al. (2024c)). However, most of these benchmarks report only a single downstream accuracy per model, without systematic means to identify the source of errors. Some further provide manually annotated error-type analyses (Zhang et al. (2024); Hao et al. (2025)), but these rely on non-standard category definitions across benchmarks, depend on post-hoc semantic inspection of reasoning traces rather than causal diagnosis, and suffer from annotator variance. In addition, the necessity of multimodal context is often left unverified (Yue et al. (2024b)). Finally, many benchmarks extract problems from publicly available sources such as textbooks, raising risks of data leakage and familiarity bias. To overcome these limitations, our study builds on FormalGeo-7K (Zhang et al. (2023)), which provides formal abstractions for both context and goal in practical mathematical geometry problems, enabling rigorous and fine-grained analysis.
@@
Approaches to adapting MLLMs for multimodal reasoning typically fall into three types: multimodal supervised finetuning, where models are trained on paired image–text inputs with ground-truth reasoning traces and answers; multimodal reinforcement learning, where models are optimized with rewards from verifiable outcomes or reasoning-trace feedback; and textual supervised finetuning, where models are tuned on large-scale textual reasoning corpora without multimodal context. These strategies are often applied sequentially or in combination, such as multimodal SFT followed by multimodal RL (e.g., Huang et al. (2025) and Yang et al. (2025)) to improve robustness, or textual SFT followed by MM-RL (e.g., Chen et al. (2025) and  Wei et al. (2025)) to transfer reasoning priors into multimodal domains. In addition, some models adopt direct RL without prior SFT (e.g., Deng et al. (2025)Meng et al. (2025)Wang et al. (2025), and Yao et al. (2025)), demonstrating that reinforcement learning alone can yield competitive reasoning performance. Closed-weight systems (e.g., OpenAI (2025)Comanici et al. (2025), and Anthropic (2025)) also report strong multimodal reasoning ability, although their training data and pipelines remain undisclosed.
@@ Appendix B Comparison with Existing Benchmarks
Multimodal mathematical problems are a popular testbed for evaluating multimodal reasoning. Among existing benchmarks (Lu et al. (2024); Wang et al. (2024b); Yue et al. (2024a)), the closest to ours is MathVerse (Zhang et al. (2024)), which, like our benchmark, primarily focuses on mathematical geometry problems and partly draws from Geometry3K (Lu et al. (2021)), a subset of FormalGeo-7K (Zhang et al. (2023)). MathVerse also provides multiple input modalities ranging from text-only to vision-only, enabling skill-specific evaluation of multimodal reasoning models, which is conceptually aligned with our design.
However, MathVerse does not meet the criteria necessary for rigorous capacity isolation: 1) it relies on curated diagrams, making it vulnerable to data leakage or familiarity effects; 2) its text descriptions are incomplete for evaluating pure reasoning, often requiring external information to be inferred from diagrams (see Figure 10); 3) overlap between question and diagram content is not explicitly controlled. Due to these limitations, analyses in this paper cannot be reproduced with the same rigor using prior benchmarks.
@@
Error type analysis offers actionable insights into model weaknesses and guides directions for improvement. However, in multimodal reasoning research, such analyses lack standardization and are almost always performed manually. Consequently, as shown in Figure 11, categories and criteria vary across datasets, making it difficult to generalize findings. Moreover, manual analysis is costly and typically applied to a single model, quickly becoming outdated along with its conclusions. By contrast, MathLens’s automatic analysis pipeline enables consistent, scalable, and up-to-date error categorization across models and datasets.
@@ C.1 MathLens
Each clause $s_{k,i}\!\in\!S_{k}$ is mapped to a corresponding set of algebraic constraints. For instance, $\textsc{Perpendicular}(AB,BC)$ is encoded as ($(x_{a}-x_{b})(x_{c}-x_{b})+(y_{a}-y_{b})(y_{c}-y_{b})=0$). Taken together, the constraints form a nonlinear system defining admissible coordinates for all points. We solve this system using the sequential least-squares programming routine (SLSQP) in scipy.optimize.minimize (Virtanen et al. (2020)), which searches for coordinates that minimize the residual of all constraints subject to feasibility bounds. To improve robustness, initial coordinates are randomly sampled, and optimization is repeated until convergence. Up to ten attempts with different random seeds are allowed, after which the sample is discarded if no feasible solution is found. Valid solutions are rendered with a matplotlib-based backend (Hunter (2007)) that draws points, line segments, arcs, and annotations according to the computed geometry. Since automatic optimization may still yield degenerate layouts (e.g., overlapping vertices, occluded labels, or extreme aspect ratios), we apply a post-processing step in which such outputs are manually filtered to preserve clarity and readability.
@@
C.2 MathLens-General
Figure 12: Data samples from MathLens-General. We curate problem instance from a prior dataset (Fu et al. (2024)) and annotate the text description and perception-centric question-answers.
Figure 13: Data samples from MathLens-General. We curate problem instance from a prior dataset (Hao et al. (2025)) and annotate the text description and perception-centric question-answers.
MathLens-General is a curated, re-annotated benchmark that extends the scope of MathLens to a broader range of images and problem domains. It includes 107 problems, each paired with an average of $\sim 7.96$ visual probe questions. Representative samples are shown in Figure 12 and Figure 13, with the data generation procedure detailed below.

### v3

@@ 3.1 Validating MathLens
To test whether MathLens captures patterns consistent with prior benchmarks, we compare it to MathVista ((Lu et al., 2024)), MathVerse ((Zhang et al., 2024)), and MathVision ((Wang et al., 2024b)) by correlating model accuracies across eight models (Figure 5, left; full results in Appendix). MathLens shows strong Spearman’s $\rho$ correlation with MathVista ($\rho=0.83$) and MathVerse ($\rho=0.86$), confirming alignment with established benchmarks, while correlation with MathVision is weaker but still positive ($\rho=0.67$). Importantly, our goal is not a new downstream benchmark but a controlled, decomposition-focused resource; high correlations therefore underscore its consistency.
@@ 3.2 How much of the gain is explained by improvement in textual reasoning?
Figure 6: Performance gains by input modality. Bars show percentage point shifts from finetuning for text versus diagram inputs. Visual gains exceed textual ones when models are primed with strong reasoning (textual SFT).
@@
Figure˜6 shows that textual SFT mainly improves textual reasoning, while multimodal RL applied afterward yields larger multimodal gains by shifting its effect to perception and integration. Direct multimodal RL without textual SFT gives only modest improvements in both modalities, and poor-quality multimodal SFT degrades multimodal performance more than textual reasoning, with RL only partially recovering the gap.
Finding 1: Multimodal RL impact varies with textual reasoning strength. With strong textual SFT it mainly boosts perception; without it, it modestly improves both modalities (Figure˜6).
@@ 3.3 How much of the gain is explained by improvement in perception?
Figure 7: Perception performance shifts from finetuning in percentage points. Except for multimodal SFT, all methods contribute to increased perception capacity.
Figure 8: Sample distribution over reflection count. Reflection is more frequent in false→true cases than in true→true cases, showing that perception gains from textual SFT arise partially from reflective reasoning.
Next, we evaluate how multimodal reasoning training affects low-level perception. For each problem $Q_{k}$, we use the perception probes $Q^{\mathrm{perc}}_{k}={q^{\mathrm{perc}}_{k,1},\ldots,q^{\mathrm{perc}}_{k,j}}$. Figure˜7 shows that all RL models contribute to better perception required for geometry problem solving. Hence, we conclude that perception is elicited even by training with correctness reward signals on the downstream problems.
Figure 10: Finetuning shifts by diagram type. Bars show percentage point changes for existing versus OOD rendered diagrams.
@@
Finding 2: Textual SFT also improves perception partially by fostering reflective reasoning (Figure˜8).
@@ 3.4 Error type analysis
MathLens’s annotation designs enable systematic categorization of model outputs into the following error categories: (1) Perception & Reasoning, failures on both perception probes and textual reasoning; (2) Perception, failure on perception probes but correct reasoning from text; (3) Reasoning, correct perception probes but failed text-based reasoning; (4) Integration, correct perception and text in isolation but failure on the combined multimodal task. Correct categories: (5) Trivial, solvable from text alone; and (6) Rest, all other correct cases.
Figure˜1 (see full results in Figure˜14 of Appendix) shows that RL primarily reduces perception and reasoning errors on the same problems, indicating correlated gains. Yet many of the reductions manifest as Integration errors, suggesting that finetuning often shifts them into failures attributed to coordination.
Finding 3: RL improves perception and reasoning in a correlated manner, and causes residual errors to be increasingly categorized as integration once the other components improve (Figure˜1).
@@ 3.5 Does multi-modal finetuning effect model’s robustness to visual inputs?
We extend the visual familiarity analysis from Section˜3.1 by testing robustness under controlled diagram variations. Instead of pixel-level augmentations (e.g., blurring), we apply semantic modifications to the geometric specification, isolating structural variation without introducing low-level artifacts (see Appendix˜C).
@@
Figure˜10 shows that while most training methods are stable under familiar diagrams, multimodal SFT suffers a sharp drop on modified diagrams. This indicates that similar downstream accuracy on public benchmarks can hide reliance on visual familiarity.
@@
To quantify robustness, we use Consistency Rate (CR) ((Zhao et al., 2024)), the expected agreement of predictions across perturbations of the same diagram:
$\Delta$ $=\mathbf{1}\!\left[M(Q^{\prime}_{k},\bar{C}_{k,i})=M(Q^{\prime}_{k},\bar{C}_{k,j})\right].$ (1)
$CR$ $=\mathbb{E}_{Q^{\prime}_{k}\sim Q^{\prime}}\Bigg[\mathbb{E}_{\begin{subarray}{c}\bar{C}_{k,i},\,\bar{C}_{k,j}\sim\bar{C}_{k}\\ i\neq j\end{subarray}}\Delta\Bigg].$ (2)
@@
Figure˜9 shows that accuracy on base diagrams is generally higher than on modified ones, indicating vulnerability to semantic perturbations. Consistency increases after multimodal RL, suggesting robustness to variation. By contrast, multimodal SFT lowers consistency.
Finding 4: Multimodal RL improves visual consistency under structural variations (Figure˜9).
@@ 3.6 Which perception skills benefit from multimodal reasoning training?
We analyze perception probe performance (Section˜3.3) by relation type, using a fine-grained categorization of geometric cues (see Section˜E.3 for full results). This analysis indicates that gains from multimodal reasoning training are not uniform across perceptual skills. Relations such as cocircular, parallel, and collinear, which depend on simple local geometric primitives, show consistent improvement. In contrast, relations such as same_angle, val_angle, triangle, and quadrilateral benefit only when textual supervision is combined with multimodal reinforcement learning, reflecting their reliance on multi-constraint reasoning and symbol–geometry alignment. By comparison, same_length, val_length, and perpendicular remain challenging, likely due to spatially displaced cues or visually ambiguous annotations. Detailed per-category results are reported in the appendix.
Finding 5: Perception gains are uneven. Multimodal reasoning training preferentially improves skills driven by direct geometric cues, while perception tasks that rely on symbolic marks or long-range visual associations show limited gains.
@@ 4 Conclusion
We introduced MathLens, a controlled benchmark that disentangles perception, reasoning, and integration in multimodal reasoning. We find that reinforcement learning primarily improves perception, especially when paired with textual supervision, while textual SFT strengthens perception indirectly through reflective reasoning. Reasoning gains largely follow perceptual improvements, leaving integration as a dominant source of residual error, and robustness diverges across strategies: RL improves consistency under diagram variation, whereas multimodal SFT reduces it due to overfitting.
Looking ahead, our results motivate future work on architectures and training strategies that more directly address cross-modal integration and perceptual capacity.
Broader Impact
Our goal is to improve the evaluation and understanding of multimodal reasoning models. By decomposing performance into perception, reasoning, and their integration, our benchmark provides more diagnostic insight than aggregate accuracy and helps clarify which subskills are responsible for observed gains.
Our work focuses on mathematical geometry problems and is constructed entirely from publicly available problem sources and synthetic derivations. It does not involve personal data, sensitive attributes, or application domains with direct societal deployment. As such, the benchmark primarily serves as a scientific analysis tool for the research community.
As with any benchmark, there is a risk that future work may over-optimize for its specific structure rather than broader generalization. To reduce this risk, the benchmark is designed with explicit subskill decompositions and controlled variants intended for diagnostic analysis rather than leaderboard-driven optimization. We do not anticipate additional ethical or societal risks beyond those commonly associated with advances in multimodal machine learning.
@@ Appendix A Related Work
Multimodal reasoning is a compositional process that requires the integration of perception and abstract reasoning ((Li et al., 2025)). Inspired by textual reasoning benchmarks, recent multimodal reasoning datasets focus on verifiable domains such as mathematics ((Lu et al., 2024; Wang et al., 2024b; Zhang et al., 2024)), scientific diagrams ((Yue et al., 2024a; Hao et al., 2025)), and charts ((Wang et al., 2024c)). However, most of these benchmarks report only a single downstream accuracy per model, without systematic means to identify the source of errors. Some further provide manually annotated error-type analyses ((Zhang et al., 2024; Hao et al., 2025)), but these rely on non-standard category definitions across benchmarks, depend on post-hoc semantic inspection of reasoning traces rather than causal diagnosis, and suffer from annotator variance. In addition, the necessity of multimodal context is often left unverified ((Yue et al., 2024b)). Finally, many benchmarks extract problems from publicly available sources such as textbooks, raising risks of data leakage and familiarity bias. To overcome these limitations, our study builds on FormalGeo-7K ((Zhang et al., 2023)), which provides formal abstractions for both context and goal in practical mathematical geometry problems, enabling rigorous and fine-grained analysis.
@@
Approaches to adapting MLLMs for multimodal reasoning typically fall into three types: multimodal supervised finetuning, where models are trained on paired image–text inputs with ground-truth reasoning traces and answers; multimodal reinforcement learning, where models are optimized with rewards from verifiable outcomes or reasoning-trace feedback; and textual supervised finetuning, where models are tuned on large-scale textual reasoning corpora without multimodal context. These strategies are often applied sequentially or in combination, such as multimodal SFT followed by multimodal RL (e.g., (Huang et al., 2025) and (Yang et al., 2025)) to improve robustness, or textual SFT followed by MM-RL (e.g., (Chen et al., 2025) and  (Wei et al., 2025)) to transfer reasoning priors into multimodal domains. In addition, some models adopt direct RL without prior SFT (e.g., (Deng et al., 2025)(Meng et al., 2025)(Wang et al., 2025), and (Yao et al., 2025)), demonstrating that reinforcement learning alone can yield competitive reasoning performance. Closed-weight systems (e.g., (OpenAI, 2025)(Comanici et al., 2025), and (Anthropic, 2025)) also report strong multimodal reasoning ability, although their training data and pipelines remain undisclosed.
@@ Appendix B Comparison with Existing Benchmarks
Multimodal mathematical problems are a popular testbed for evaluating multimodal reasoning. Among existing benchmarks ((Lu et al., 2024; Wang et al., 2024b; Yue et al., 2024a)), the closest to ours is MathVerse ((Zhang et al., 2024)), which, like our benchmark, primarily focuses on mathematical geometry problems and partly draws from Geometry3K ((Lu et al., 2021)), a subset of FormalGeo-7K ((Zhang et al., 2023)). MathVerse also provides multiple input modalities ranging from text-only to vision-only, enabling skill-specific evaluation of multimodal reasoning models, which is conceptually aligned with our design.
However, MathVerse does not meet the criteria necessary for rigorous capacity isolation: 1) it relies on curated diagrams, making it vulnerable to data leakage or familiarity effects; 2) its text descriptions are incomplete for evaluating pure reasoning, often requiring external information to be inferred from diagrams (see Figure˜11); 3) overlap between question and diagram content is not explicitly controlled. Due to these limitations, analyses in this paper cannot be reproduced with the same rigor using prior benchmarks.
@@
Error type analysis offers actionable insights into model weaknesses and guides directions for improvement. However, in multimodal reasoning research, such analyses lack standardization and are almost always performed manually. Consequently, as shown in Figure˜12, categories and criteria vary across datasets, making it difficult to generalize findings. Moreover, manual analysis is costly and typically applied to a single model, quickly becoming outdated along with its conclusions. By contrast, MathLens’s automatic analysis pipeline enables consistent, scalable, and up-to-date error categorization across models and datasets.
@@ C.1 MathLens
Each clause $s_{k,i}\!\in\!S_{k}$ is mapped to a corresponding set of algebraic constraints. For instance, $\textsc{Perpendicular}(AB,BC)$ is encoded as ($(x_{a}-x_{b})(x_{c}-x_{b})+(y_{a}-y_{b})(y_{c}-y_{b})=0$). Taken together, the constraints form a nonlinear system defining admissible coordinates for all points. We solve this system using the sequential least-squares programming routine (SLSQP) in scipy.optimize.minimize ((Virtanen et al., 2020)), which searches for coordinates that minimize the residual of all constraints subject to feasibility bounds. To improve robustness, initial coordinates are randomly sampled, and optimization is repeated until convergence. Up to ten attempts with different random seeds are allowed, after which the sample is discarded if no feasible solution is found. Valid solutions are rendered with a matplotlib-based backend ((Hunter, 2007)) that draws points, line segments, arcs, and annotations according to the computed geometry. Since automatic optimization may still yield degenerate layouts (e.g., overlapping vertices, occluded labels, or extreme aspect ratios), we apply a post-processing step in which such outputs are manually filtered to preserve clarity and readability.
@@
C.2 MathLens-NS
MathLens-NS is a curated, re-annotated benchmark that extends the scope of MathLens to a broader range of images and problem domains. It includes 107 problems, each paired with an average of $\sim 7.96$ visual probe questions. Representative samples are shown in Figure˜15 and Figure˜16, with the data generation procedure detailed below.